### v6

@@ Abstract
Neural electrical activity is fundamental to brain function, underlying a range of cognitive and behavioral processes, including movement, perception, decision-making, and consciousness. Abnormal patterns of neural signaling often indicate the presence of underlying brain diseases. The variability among individuals, the diverse array of clinical symptoms from various brain disorders, and the limited availability of diagnostic classifications, have posed significant barriers to formulating reliable model of neural signals for diverse application contexts. Here, we present BrainWave, the first foundation model for both invasive and non-invasive neural recordings, pretrained on more than 40,000 hours of electrical brain recordings (13.79 TB of data) from approximately 16,000 individuals. Our analysis show that BrainWave outperforms all other competing models and consistently achieves state-of-the-art performance in the diagnosis and identification of neurological disorders. We also demonstrate robust capabilities of BrainWave in enabling zero-shot transfer learning across varying recording conditions and brain diseases, as well as few-shot classification without fine-tuning, suggesting that BrainWave learns highly generalizable representations of neural signals. We hence believe that open-sourcing BrainWave will facilitate a wide range of clinical applications in medicine, paving the way for AI-driven approaches to investigate brain disorders and advance neuroscience research.
@@ 1 Introduction
Electrical brain recordings, capturing the intricate patterns of the brain’s electrical activities, are essential in advancing our understanding of brain across scientific domains [1, 2, 3, 4, 5, 6, 7]. In particular, they can be used to identify medical conditions and diagnose neurological disorders, and are thus essential for addressing major global health challenges in developing nations [8, 9]. Scalp electroencephalography (EEG) and intracranial electroencephalography (iEEG) are two primary methods to conduct these types of recordings. EEG is non-invasive and economical viable, and has thus been used in diverse applications [10, 11, 12, 13, 14]. In contrast, iEEG offers high signal fidelity and spatial resolution [15], but the invasive nature limits its applicability to only the most severe patient cases and restricted scenarios. Due to the distinct features of EEG and iEEG data, such as different acquisition rates and notable variations in channel numbers [16, 17, 18], studies have so far investigated them separately. We hypothesize that combining EEG and iEEG data can offer information that are not only rich in detail but also highly generalizable across diverse neural electrical activities, and develop a pioneering foundational model, BrainWave, for both EEG and iEEG data. BrainWave learns robust representations that achieve the state-of-art performance in a wide range of tasks, demonstrating the synergy of EEG and iEEG data for the first time.
@@
BrainWave overcomes a suite of inherent challenges in conventional supervised artificial intelligence (AI) models used for the analysis of brain signals. First, BrainWave leverages self-supervised training, circumventing the need for large-scale, high-quality manual labeling. In clinical applications, the process of brain data annotation is labor-intensive and reliant on specialized expertise [19, 20, 21], exemplified by the need for multi-day monitoring for epilepsy patients [22] and the clinical experts’ capacity to annotate only tens of seconds of data in a single work period. Second, BrainWave provides much-need generalization at two levels that were not possible in supervised training, which requires an understanding of fundamental and general patterns in brain signals. Individual variability in brain neural activities, a consequence of each person’s distinct brain structure and functional behaviors [23], leads to markedly different brain recording patterns [24]. This diversity hinders the generalizability of supervised AI models, as they often struggle to extend the insights gained from a subset of patients to a broader population, due to significant differences across individuals, as well as variability that changes with different behavioral states. Moreover, there are numerous types of brain-related diseases with various underlying mechanisms  [25, 26, 27, 28], and even a single disease may present with multiple subtypes [29]. Supervised AI models are also task-specific and fail to address a diverse range of tasks using brain signals [30, 31, 32, 24, 33, 34, 35, 36, 37], indicating that they do not actually “understand” brain signals.
@@
While prior work has attempted to build foundation models for brain signals [38, 39, 40, 41, 42, 43], BrainWave offered unique technical contributions by building the largest dataset of electrical brain recordings and developing novel techniques to integrate EEG and iEEG data for the first time. As illustrated in Figure 1a, we collected a total of 13.79 TB of combined EEG and iEEG data over a duration of 40,907 hours, which serves as the foundation for pretraining BrainWave. The data were obtained from 15,997 individuals, including both healthy individuals and those with various brain disorders, spanning an age range from infancy ($<1$ year) to over 90 years old. In the pretraining stage, we adopt a masked modeling strategy that reconstructs the time-frequency representations of the masked patches (Fig. 1b). Unlike previous works that uniformly resampled data to a common sampling rate [44, 45, 46], the designed embedding layer of BrainWave can adapt to different temporal resolutions, which enhanced the data scalability and provided the model with stronger generalization capabilities. We also employed a channel count-agnostic approach to capture the inter-channel relationships. Empowered by growing datasets and advances in model design, BrainWave exhibited highly robust pretrained representations and excellent transfer learning capabilities (Fig. 1c).
@@
Figure 1: Overview of BrainWave. a, Data curation for pretraining BrainWave. The pretraining corpus contains both invasive and non-invasive brain recordings collected from diverse healthcare scenarios. b, The pretraining pipeline of BrainWave. BrainWave is pretrained on more than 3 billion signal patches using a masked modeling strategy. c, The evaluation tasks consist of few-shot classification and cross-domain evaluation. We conduct few-shot classification with a prototypical network in which the we directly compare the representations of the queries with class prototypes. We perform three different levels of cross-domain analysis: cross-subject, cross-hospital, and cross-subtype. d, The overall results of BrainWave compared to other pretrained models. BrainWave outperforms other models across all the 28 experiments, with significant improvement ($p<0.001$) in 24 of them.
@@
We systematically evaluated BrainWave and observed its exceptional performance across different settings, including cross-subject (Fig. 2a and b), cross-hospital (Fig. 2c) and cross-subtype (Fig. 2d) tasks, which attests to its strong generalization capabilities. Additionally, its proficiency in few-shot classification (Fig. 3) highlights its adaptability in learning from limited data. We also demonstrated the effectiveness of joint pretraining by leveraging both invasive and non-invasive neural data (Fig. 4a-c) and found that it achieved superior performance by learning more generalizable representations (Fig. 4e) and richer semantic information (Supplementary Tables 38 and 39). To comprehensively assess the capabilities of BrainWave as a foundation model for electrical brain recordings in healthcare scenarios, we constructed a benchmark comprising 15 different datasets. BrainWave is compared against the previous state-of-the-art foundation models that are publicly available, including LaBraM [44], BrainBERT [46] and MOMENT [47]. Figure 1d summarizes the overall results of BrainWave compared with other methods, in which BrainWave attains consistently state-of-the-art performance on all the 28 experiments, with significant improvement ($p<0.001$) over the second-best method in 24 experiments, showing the versatility of BrainWave in a wide array of tasks for brain recordings in healthcare. To investigate the effectiveness of joint pretraining with both invasive and non-invasive neural data, we separately pretrained two model variants using EEG and iEEG data under the condition of maintaining other pretraining settings constant, and compared them across all experimental settings (Fig. 4a-d). The results show that BrainWave outperforms other variants in diverse tasks and experimental setups, marking it as the first to successfully implement joint pretraining and validate its effectiveness. Overall, we demonstrated the potential of BrainWave to assist clinical diagnostics and decision support with strong scalability.
@@
BrainWave not only exhibits potential for broad applications but also offers new insights for the development of foundation models based on brain signals. We will release BrainWave as an off-the-shelf and publicly available model, which will serve as a basis for others in their own tasks, facilitating diverse clinical applications and research for brain recordings.
@@ 2 Results
Figure 2: Performance of cross-domain evaluation. a-l, Bar plots comparing the AUROC scores of BrainWave and competing models on cross-subject tasks. Each experiment is conducted with $n$-fold cross validation ($n$ is the number of subject groups), where we repeat five runs for each fold. m,n, Bar plots comparing the AUROC scores of BrainWave and competing models on cross-hospital tasks. o,p, Bar plots comparing the AUROC scores of BrainWave and competing models on cross-subtype tasks. m-p, The source dataset is served as the training set and the target dataset is served as the evaluation set. In each experiment, we repeat five runs. a-p, Data are mean $\pm$ SD. The listed $p$ value indicates the significance for BrainWave outperforming the best comparison model, with the two-sided $t$-test.
@@ 2.1 Cross-domain Disease Diagnosis and Detection
In real-world clinical applications, where the data of individuals to be diagnosed are usually unavailable during training, it is desirable for a model to be capable of making accurate predictions on unseen subjects. Thus, evaluating in a cross-subject setting is essential to accurately reflect a model’s performance and practical value in clinical scenarios. For each dataset, we split the subjects into $n$ non-overlapping groups and employed $n$-fold cross validation to evaluate all models, ensuring that all subjects are included in the testing process. In each fold, we randomly selected a subject group from the training set for validation and repeated five runs. BrainWave was evaluated against competing methods on a total of 12 datasets, consisting of 10 publicly available datasets (Supplementary Tables 7-16) and 2 private datasets (Supplementary Tables 3 and 4). Model performance was reported using the area under the receiver operating curve (AUROC) and balanced accuracy (BACC). We calculated $p$ values with the two-sided $t-$test between BrainWave and the most competitive comparison model for each task to check for significance.
@@
Across all experiments, BrainWave consistently outperformed other methods with an average relative improvement of 12.61% in AUROC (Area Under the Receiver Operating Characteristic curve) and 16.44% in BACC (Balanced Accuracy) over each second-best performing model (Fig. 2a-l and Extended Data Fig. 5). In schizophrenia diagnosis (Fig. 2i; dataset Schizophrenia-28 [48]), BrainWave achieved an improvement of 37.44% and 41.59% compared to the best comparison model in terms of AUROC and BACC. In seizure detection (Fig. 2b; dataset CHB-MIT [49]), BrainWave demonstrated a 26.18% boost relative to the second best method. The strong performance of BrainWave on both ADHD-Adult [50] (Fig. 2j) and ADHD-Child [51] (Fig. 2k) indicated its robustness across age groups, owing to the broad age distribution in our pretraining corpus. In conclusion, BrainWave significantly surpassed other models ($p<0.001$) on 11 out of the 12 datasets, demonstrating its superiority in disease diagnosis and detection.
@@
The promising results of BrainWave on cross-subject evaluation motivated us to further explore its transfer capability across more divergent distributions. Thus, we attempted two more challenging experimental setups. Under these settings, we fine-tuned on one dataset and directly apply the model to another. These datasets are not only collected from different individuals but also from different hospitals and collection devices, or from patients with different disease subtypes. Unlike the traditional paradigm where fine-tuning is dataset-specific, these settings allow the model to be seamlessly deployed across different datasets and even different but related tasks. The cross-hospital evaluation involved mutual transfer between two datasets, Mayo-Clinic and FNUSA [52], both collected from patients with drug resistant epilepsy but from different hospitals. The cross-subtype evaluation included three datasets, namely Absence-16, Clonic-6, and Atonic-5, collected from patients with different subtypes of seizures (absence seizure, clonic seizure, and atonic seizure), and we performed zero-shot transfer from Absence-16 to Clonic-6 and Atonic-5. BrainWave showed promising results and achieved the best performance in all cross evaluations (Fig. 2m-p and Extended Data Fig. 6). For instance, BrainWave achieved an impressive AUROC of 93.82% in the zero-shot transfer from FNUSA to Mayo-Clinic (Fig. 2n). When transferring across different seizure subtypes (Fig. 2o,p), BrainWave exhibited average improvements of 13.90% and 13.49% in terms of AUROC and BACC, respectively, compared to the second-best model.
In summary, our experimental results indicated that BrainWave holds the potential to reduce labeling and training costs under similar scenarios, demonstrating promising applicability in real-world clinical tasks.
@@
Figure 3: Performance and analysis of few-shot classification. a-l, Box plots comparing the AUROC scores of BrainWave and competing models on few-shot classification. We conduct $n$-fold cross validation for each experiment and repeat five runs per fold. We perform 3-shot and 8-shot classification for each task. m, Box plots comparing the BACC scores of BrainWave on few-shot classification and end-to-end trained MLP with full-label supervision. n, t-SNE (t-distributed Stochastic Neighbor Embedding) plots of the pretrained representations on Absence-16 generated from BrainWave and other pretrained encoders. Each model contains four subplots, with each subplot generated by randomly sampling a portion of the original dataset.
@@ 2.2 Few-shot Classification
In clinical practice, limited availability of labeled data sometimes poses challenges for fine-tuning models, which highlights the critical importance of learning sufficiently robust representations. Consequently, we performed few-shot classification, which is an evaluation scheme that studies the generalization capabilities of models on new tasks given a very limited number of labeled examples. We adopted a direct comparison strategy for classification by comparing the representations of the queries with prototype of each category (Fig. 1c). The process solely involved obtaining representations from the pretrained models and computing class prototypes, without any parameter updates or introduction of new parameters. The few-shot experiments were still conducted under the cross-subject setting and we employed $n$-fold cross validation, where in each fold we randomly chose labeled examples from the training set as the support set. We established two sizes for the support set, with 3 and 8 labeled examples per class (3-shot and 8-shot), respectively. Given that the performance can fluctuate depending on the support set, we repeated experiments over five runs in each fold. For comparison, we also trained an MLP (multilayer perceptron) from scratch with full-label supervision for each experiment.
@@
We conducted few-shot experiments (Fig. 3a-l and Extended Data Fig. 7) on all datasets used in the cross-subject evaluation and found that BrainWave still maintained solid performance by achieving an average improvement of 21.21% in terms of AUROC compared to the second-best model. For instance, on Absence-16 (Fig. 3c) and ADHD-Adult (Fig. 3j), BrainWave achieved an AUROC over 90% (91.93% on Absence-16, 90.39% on ADHD-Adult) in 8-shot classification. On MDD-64 [53] (Fig. 3g), the performance of 8-shot learning even nearly matched that of full-label supervised fine-tuning (89.83% versus 91.50%). BrainWave not only outperformed other models by a large margin but also demonstrated greater robustness to the selection of the support set. Specifically, we calculated the average standard deviation across all few-shot experiments and found that the fluctuations in BrainWave’s performance are, on average, smaller than those of the second-best performing models (5.74% versus 6.06%). The poor performance of BrainBERT in few-shot classification may be due to its limitation in supporting only fixed-length inputs, which requires additional operations to accommodate varying input lengths, highlighting the importance of supporting variable-length inputs for a wide array of tasks.
@@
When compared to the end-to-end trained MLP (Fig. 3m), BrainWave still achieved an average improvement of 25.79% and 26.60% in terms of AUROC and BACC, respectively. Surprisingly, we also observed that when comparing the 8-shot performance of BrainWave with the full-label fine-tuning (i.e., with thousands or even tens of thousands of labeled examples) performance of other pretrained models, our model still outperforms on the majority of datasets (Extended Data Figs. 9 and 10). Specifically, BrainWave surpassed LaBraM, BrainBERT and MOMENT on 7, 12 and 11 out of 12 datasets, respectively. To better illustrate the results, we visualized the representations of the pretrained models, with different colored points representing different categories (Fig. 3n) and Extended Data Fig. 11). Even without fine-tuning, the representations generated by BrainWave are sufficiently discriminative and enable its outstanding performance in few-shot classification. Overall, our extensive evaluation of few-shot classification demonstrated the immense potential of BrainWave as a foundational model for electrical brain recordings that offers robust and off-the-shelf representations for various clinical applications.

### v7

@@ Abstract
Neural electrical activity is fundamental to brain function, and abnormal patterns of neural signaling often indicate the presence of underlying brain diseases. The variability among individuals, the diverse array of clinical symptoms from various brain disorders, and the limited availability of diagnostic classifications, have posed significant barriers to formulating reliable models of neural signals for diverse application contexts. Here, we present BrainWave, the first foundation model for both invasive and non-invasive neural recordings, pretrained on more than 40,000 hours of electrical brain recordings (13.79 TB of data) from approximately 16,000 individuals. Our analysis show that BrainWave consistently achieves state-of-the-art performance in the identification of neurological disorders across various experimental settings. In addition, we demonstrate the effectiveness of pretraining, as BrainWave achieves strong few-shot classification performance without fine-tuning, indicating our pretraining strategy extracts informative representations from neural signals. BrainWave is also evaluated in real-world clinical scenarios, highlighting its potential in facilitating clinical interpretation and decision-making. We hence believe that open-sourcing BrainWave will facilitate a wide range of clinical applications in medicine, paving the way for AI-driven approaches to investigate brain disorders.
@@ 1 Introduction
Electrical brain recordings, capturing the intricate patterns of the brain’s electrical activities, are essential in advancing our understanding of brain across scientific domains [1, 2, 3, 4, 5, 6, 7]. In particular, they can be used to identify medical conditions and diagnose neurological disorders, and are thus essential for addressing major global health challenges in developing nations [8, 9]. Scalp electroencephalography (EEG) and intracranial electroencephalography (iEEG) are two primary methods to conduct these types of recordings. EEG is non-invasive and economical viable, and has thus been used in diverse applications [10, 11, 12, 13, 14]. In contrast, iEEG offers high signal fidelity and spatial resolution [15], but the invasive nature limits its applicability to only the most severe patient cases and restricted scenarios. Due to the distinct features of EEG and iEEG data, such as different acquisition rates and notable variations in channel numbers [16, 17, 18], studies have so far investigated them separately. We hypothesize that combining EEG and iEEG data can offer information that are not only rich in detail but also highly generalizable across diverse neural electrical activities, and develop a pioneering foundational model, BrainWave, for both EEG and iEEG data. BrainWave learns robust representations that achieve state-of-the-art performance in a wide range of tasks, demonstrating the synergy of EEG and iEEG data for the first time.
@@
BrainWave overcomes a suite of inherent challenges in conventional supervised artificial intelligence (AI) models used for the analysis of brain signals. First, BrainWave leverages self-supervised training, circumventing the need for large-scale, high-quality manual labeling. In clinical applications, the process of brain data annotation is labor-intensive and reliant on specialized expertise [19, 20, 21], exemplified by the need for multi-day monitoring for epilepsy patients [22] and the clinical experts’ capacity to annotate only tens of seconds of data in a single work period. Second, BrainWave provides much-need generalization at two levels that were not possible in supervised training, which requires an understanding of fundamental and general patterns in brain signals. Individual variability in brain neural activities, a consequence of each person’s distinct brain structure and functional behaviors [23], leads to markedly different brain recording patterns [24]. This diversity hinders the generalizability of supervised AI models, as they often struggle to extend the insights gained from a subset of patients to a broader population, due to significant differences across individuals, as well as variability that changes with different behavioral states. Moreover, there are numerous types of brain-related diseases with various underlying mechanisms  [25, 26, 27, 28], and even a single disease may present with multiple subtypes [29]. Supervised AI models are task-specific and fail to address a diverse range of tasks using brain signals [30, 31, 32, 24, 33, 34, 35, 36, 37].
@@
While prior work has attempted to build foundation models for brain signals [38, 39, 40, 41, 42, 43], BrainWave offered unique technical contributions by building the largest dataset of electrical brain recordings and developing novel techniques to integrate EEG and iEEG data for the first time. As illustrated in Figure 1a, we collected a total of 13.79 TB of combined EEG and iEEG data over a duration of 40,907 hours, which serves as the foundation for pretraining BrainWave. The data were obtained from 15,997 individuals, including both healthy individuals and those with various brain disorders, spanning an age range from infancy (<1<1 year) to over 90 years old. The sampling rates vary across different brain signal datasets, with a more pronounced disparity between EEG and iEEG data, which increases the difficulty of unified modeling. Previous works generally focused on modeling either EEG or iEEG alone, so they often uniformly resampled the data to a common sampling rate [44, 45, 46]. We proposed a scale alignment layer that adapts to arbitrary temporal resolutions without the need for resampling, thereby improving data scalability and enhancing generalization capabilities. In the pretraining stage, we adopt a masked modeling strategy that reconstructs the time-frequency representations of the masked patches (Fig. 1b). We also employed a channel count-agnostic approach to capture the inter-channel relationships. Empowered by growing datasets and advances in model design, BrainWave exhibited highly robust pretrained representations and excellent transfer learning capabilities (Fig. 1c).
@@
Figure 1: Overview of BrainWave. a, Data curation for pretraining BrainWave. The pretraining corpus contains both invasive and non-invasive brain recordings collected from diverse healthcare scenarios. b, The pretraining pipeline of BrainWave. BrainWave is pretrained on more than 3 billion signal patches using a masked modeling strategy. c, The evaluation tasks consist of few-shot classification and cross-domain evaluation. We conduct few-shot classification with a prototypical network in which the we directly compare the representations of the queries with class prototypes. We perform three different levels of cross-domain analysis: cross-subject, cross-hospital, and cross-subtype. d, The overall results of BrainWave compared to other pretrained models. BrainWave outperforms other models across all the 24 experiments, with significant improvement ($p<0.001$) in 20 of them.
@@
To comprehensively assess the capabilities of BrainWave as a foundation model for electrical brain recordings in healthcare scenarios, we constructed a benchmark consisting of 15 datasets and 20 tasks. To evaluate generalization, we systematically assessed BrainWave across different cross-domain settings, including cross-subject (Fig. 2a and b), cross-hospital (Fig. 2c) and cross-subtype (Fig. 2d) tasks. To verify the effectiveness of pretraining, we introduced few-shot classification tasks to evaluate the capability of the pretrained representations when directly applied to downstream tasks without fine-tuning (Fig. 3). BrainWave is compared against the previous state-of-the-art foundation models that are publicly available, including LaBraM [44], BrainBERT [46] and MOMENT [47]. Figure 1d summarizes the overall results of BrainWave compared with other methods, in which BrainWave attains consistently state-of-the-art performance on all the 24 experiments, with significant improvement ($p<0.001$) over the second-best method in 20 experiments, showing the versatility of BrainWave in a wide array of tasks. To demonstrate the practical value of BrainWave, we designed a series of clinical tasks in two real-world scenarios: seizure onset zone localization in epilepsy, and prediction of key biomarkers and clinical scale scores in Alzheimer’s disease, showcasing its potential in supporting clinical decision-making (Fig. 5). To investigate the effectiveness of joint pretraining with both non-invasive and invasive neural data, we compared BrainWave with two model variants pretrained exclusively on EEG or iEEG data across all experiments (Fig. 5a-d). The results show that BrainWave outperforms other variants in diverse tasks, highlighting the effectiveness of its design empowered by joint pretraining. In conclusion, our study demonstrates (1) the robustness of the pretrained representations, (2) the potential of BrainWave in clinical diagnostic support, and (3) the effectiveness of joint pretraining.
@@
We will release BrainWave as a publicly available model, which will serve as a basis for others in their own tasks, facilitating diverse clinical applications and research for brain recordings.
@@ 2 Results
Figure 2: Performance of cross-domain evaluation. a-j, Bar plots comparing the AUROC scores of BrainWave and competing models on cross-subject tasks. Each experiment is conducted with $n$-fold cross validation ($n$ is the number of subject groups), where we repeat five runs for each fold. k,l, Bar plots comparing the AUROC scores of BrainWave and competing models on cross-hospital tasks. m,n, Bar plots comparing the AUROC scores of BrainWave and competing models on cross-subtype tasks. k-n, The source dataset is served as the training set and the target dataset is served as the evaluation set. In each experiment, we repeat five runs. a-n, Data are mean $\pm$ SD. The listed $p$ value indicates the significance for BrainWave outperforming the best comparison model, with the two-sided $t$-test.
@@ 2.1 Cross-domain Disease Diagnosis and Detection
Cross-domain evaluation is an important setting for verifying the generalization capability of the model. Firstly, we conducted cross-subject evaluation, in which we split the subjects into $n$ non-overlapping groups and employed $n$-fold cross validation to evaluate all models, ensuring that all subjects are included in the testing process. In each fold, we randomly selected a subject group from the training set for validation and repeated five runs. BrainWave was evaluated against competing methods on a total of 10 datasets, including Alzheimer’s disease (AD), epilepsy, major depressive disorder, schizophrenia and attention-deficit/hyperactivity disorder (ADHD). Model performance was reported using the area under the receiver operating curve (AUROC) and balanced accuracy (BACC). We calculated $p$ values with the two-sided $t$-test between BrainWave and the most competitive comparison model for each task to check for significance. Across all experiments, BrainWave consistently outperformed other methods with an average relative improvement of 11.93% in AUROC and 17.59% in BACC over each second-best performing model (Fig. 2a-j and Extended Data Fig. 6). In schizophrenia diagnosis (Fig. 2h; dataset Schizophrenia-28 [48]), BrainWave achieved an improvement of 37.44% and 41.59% compared to the best comparison model in terms of AUROC and BACC. In seizure detection (Fig. 2b; dataset CHB-MIT [49]), BrainWave demonstrated a 26.18% boost relative to the second best method. The strong performance of BrainWave on both ADHD-Adult [50] (Fig. 2i) and ADHD-Child [51] (Fig. 2j) indicated its robustness across age groups, owing to the broad age distribution in our pretraining corpus. In conclusion, BrainWave significantly surpassed other models ($p<0.001$) on 9 out of the 10 datasets, demonstrating its superiority in disease diagnosis and detection.
@@
The promising results of BrainWave on cross-subject evaluation motivated us to further explore its transfer capability across more divergent distributions. Thus, we attempted two more challenging experimental setups. Under these settings, we fine-tuned on one dataset and directly apply the model to another. These datasets are not only collected from different individuals but also from different hospitals and collection devices, or from patients with different disease subtypes. Unlike the traditional paradigm where fine-tuning is dataset-specific, these settings allow the model to be seamlessly deployed across different datasets and even different but related tasks. The cross-hospital evaluation involved mutual transfer between two datasets, Mayo-Clinic and FNUSA [52], both collected from patients with drug resistant epilepsy (DRE) but from different hospitals. The cross-subtype evaluation included three datasets, namely Absence-16, Clonic-6, and Atonic-5, collected from patients with different subtypes of seizures (absence seizure, clonic seizure, and atonic seizure), and we performed zero-shot transfer from Absence-16 to Clonic-6 and Atonic-5. BrainWave showed promising results and achieved the best performance in all cross evaluations (Fig. 2m-p and Extended Data Fig. 7). For instance, BrainWave achieved an impressive AUROC of 93.82% in the zero-shot transfer from FNUSA to Mayo-Clinic (Fig. 2n). When transferring across different seizure subtypes (Fig. 2o,p), BrainWave exhibited average improvements of 13.90% and 13.49% in terms of AUROC and BACC, respectively, compared to the second-best model. In summary, our experimental results indicated that BrainWave holds the potential to reduce labeling and training costs under similar scenarios, demonstrating strong transferability across disease detection tasks.
@@
Figure 3: Performance and analysis of few-shot classification. a-j, Box plots comparing the AUROC scores of BrainWave and competing models on few-shot classification. We conduct $n$-fold cross validation for each experiment and repeat five runs per fold. We perform 3-shot and 8-shot classification for each task. k, t-SNE (t-distributed Stochastic Neighbor Embedding) plots of the pretrained representations on Absence-16 generated from BrainWave and other pretrained encoders. Each model contains four subplots, with each subplot generated by randomly sampling a portion of the original dataset.
@@ 2.2 Few-shot Classification
In clinical practice, limited availability of labeled data sometimes poses challenges for fine-tuning models, which highlights the critical importance of learning sufficiently robust representations. Consequently, we performed few-shot classification, which is an evaluation scheme that studies the generalization capabilities of models on new tasks given a very limited number of labeled examples. We adopted a direct comparison strategy for classification by comparing the representations of the queries with prototype of each category (Fig. 1c). The process solely involved obtaining representations from the pretrained models and computing class prototypes, without any parameter updates or introduction of new parameters. The few-shot experiments were still conducted under the cross-subject setting and we employed $n$-fold cross validation, where in each fold we randomly chose labeled examples from the training set as the support set. We established two sizes for the support set, with 3 and 8 labeled examples per class (3-shot and 8-shot), respectively. Given that the performance can fluctuate depending on the support set, we repeated experiments over five runs in each fold.
@@
We conducted few-shot experiments (Fig. 3a-j and Extended Data Fig. 8) on all datasets used in the cross-subject evaluation and found that BrainWave still maintained solid performance by achieving an average improvement of 22.23% in terms of AUROC compared to the second-best model. For instance, on Absence-16 (Fig. 3c) and ADHD-Adult (Fig. 3i), BrainWave achieved an AUROC over 90% (91.93% on Absence-16, 90.39% on ADHD-Adult) in 8-shot classification. On MDD-64 [53] (Fig. 3f), the performance of 8-shot learning even nearly matched that of full-label supervised fine-tuning (89.83% versus 91.50%). BrainWave not only outperformed other models by a large margin but also demonstrated greater robustness to the selection of the support set. Specifically, we calculated the average standard deviation across all few-shot experiments and found that the fluctuations in BrainWave’s performance are, on average, smaller than those of the second-best performing models (5.74% versus 6.06%). The poor performance of BrainBERT in few-shot classification may be due to its limitation in supporting only fixed input length and sampling rate. This limitation requires additional operations to accommodate varying datasets, highlighting the importance of flexible input support for diverse tasks.
@@
Surprisingly, we also observed that when comparing the 8-shot performance of BrainWave with the full-label fine-tuning (i.e., with thousands or even tens of thousands of labeled examples) performance of other pretrained models, our model still outperforms on the majority of datasets (Extended Data Fig. 9). Across all datasets, BrainWave achieved average AUROC improvements of 2.27%, 26.40% and 14.87% over LaBraM, BrainBERT and MOMENT, respectively. To better illustrate the results, we visualized the representations of the pretrained models, with different colored points representing different categories (Fig. 3k) and Extended Data Fig. 10). Even without fine-tuning, the representations generated by BrainWave are sufficiently discriminative and enable its outstanding performance in few-shot classification. Overall, our extensive evaluation of few-shot classification demonstrated the immense potential of BrainWave as a foundational model that offers robust and off-the-shelf representations for electrical brain recordings.
Figure 4: Clinical tasks on epilepsy and AD. a, Channel-level epileptic waveform detection on 4 patients with DRE. Sensitivity/Specificity and AUROC/AP are reported. b, The process of pinpointing channels with frequent epileptic discharges and repeated involvement as seizure onset sites. We can quickly quantify these metrics for each channel from model predictions. c-g, Predictions of amyloid-beta deposition and a series of clinical scale scores from AD patients. Each experiment is conducted with 5-fold cross validation, where we repeat five runs for each fold. c, Amyloid-beta deposition prediction. d, MMSE score prediction. We divide it into 4 discrete ranges: 24-30, 21-23, 10-20 and 0-9. e, MoCA-B score prediction. We divide it into 4 discrete ranges: 26-30, 18-25, 10-17 and 0-9. f, ROCF score prediction. We divide it into 5 discrete ranges: 33-36, 24-32, 18-23, 12-17 and 0-11. g, PSQI score prediction. We divide it into 4 discrete ranges: 0-5, 6-10, 11-15 and 16-21.
2.3 Clinical Application
In addition to its superior performance on brain disorder detection benchmarks, BrainWave demonstrates practical utility by offering valuable assistance and guidance in real-world clinical scenarios.
Firstly, we utilized BrainWave to assist in localizing the seizure onset zone (SOZ), which refers to the region of the brain where epileptic seizures originate. Accurate identification of the SOZ is critical for planning surgical treatment in patients with DRE, as it guides the localization of the epileptogenic focus and supports diagnostic evaluation for subsequent resective surgery. Patients with DRE often require intracranial electrode implantation to record seizure activity. The SOZ can be localized by analyzing the waveforms of different channels, which involves distinguishing seizure patterns on a per-channel basis. Such a process demands more fine-grained and precise analysis than merely classifying epileptic versus normal waveforms over a certain time period. As shown in Fig. 4a, BrainWave is evaluated on channel-level epileptic waveform detection across 4 patients with DRE, in which the average Sensitivity/Specificity, AUROC/Average Precision (AP) achieved 84.65%/86.87% and 93.24%/86.71% respectively. The predictions of BrainWave helped reveal the spatial distribution of seizure activity. We took patient 04 as an example. The heatmap shown on the left side of Fig 4b is the channel-level predicted probabilities of multiple seizures from patient 04 provided by BrainWave. As shown in the middle of Fig 4b, the predictions enabled us to quantify two metrics for each channel: the probability of epileptic discharge occurrence across multiple seizures (upper), and the number of times it serves as the seizure onset site (lower). Higher values of these two metrics suggested a greater likelihood of the channel being part of the SOZ. As shown on the right side of Fig 4b, clinicians can quickly identify the corresponding brain regions based on the statistical results, facilitating surgical planning.
Secondly, BrainWave can predict clinical assessment indicators for AD based on EEG signals. The diagnosis of AD involves a series of physiological examinations and cognitive scale tests. The results of these assessments serve as important references for the clinical diagnosis of AD. We conducted experiments on 6 patients, using their recorded EEG data to predict amyloid-beta deposition and a range of clinical scale scores (Supplementary Tables 17), including Mini-Mental State Examination (MMSE), Montreal Cognitive Assessment-Basic (MoCA-B), Rey–Osterrieth Complex Figure Test (ROCF), and Pittsburgh Sleep Quality Index (PSQI). Amyloid-beta deposition is one of the key biomarkers in the diagnosis of AD, which is often detected by Positron Emission Tomography (PET) imaging. MMSE, MoCA-B, and ROCF are cognitive assessment tests, and PSQI is used to evaluate sleep quality. For the clinical scale scores, we divided the scores into several discrete ranges for prediction. Each range represents an evaluation category, such as normal, mild cognitive impairment, moderate cognitive impairment, severe cognitive impairment, and so on. We split the data into 5 groups and utilized a 5-fold cross validation. Each fold was repeated 5 runs. Fig 4c-g shows the results, in which BrainWave achieved an average accuracy close to or exceeding 95% across all tasks, along with an average Cohen’s Kappa above 0.9. BrainWave’s predictions exhibited strong consistency with established clinical evaluation outcomes, highlighting its great potential to support clinical diagnosis.